if we join the bases of 2 identical cylindrical tubes. the height of each tube is 15 cm and the diameter is 6 cm. what will be the surface area of new shape?
Answer by addingup(3677)

15+15 = h (height of new cylinder)
6*Pi = perimeter of the circumference (the tube)
(6*Pi)*(15+15)= area of the tube
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Note: surface area of the open tube, without caps at the ends. If you want caps at the ends, to the result you get you have to add:(6/2)^2*Pi, or you can do it this way:(Pi/4)*6^2
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Practice using your calculator.
J